OpenStudy (anonymous):

WHAT IS A QUADRATIC EQUATION ... HOW CAN I SOLVE IT .... PLEASE HELp.... GIVE RELEVANT FORMULAS .... I AM STUCK :'(

OpenStudy (anonymous):

i didnt get it !! new to algebra ... how can u use alphabets in maths ???

OpenStudy (anonymous):

maths has something to do with numbers :'( .... help

Parth (parthkohli):

The great Aristotle defined mathematics as "the science of quantity".

OpenStudy (anonymous):

^^ see .... can anyone explain the use of alphabets in maths ????

Parth (parthkohli):

Letters are used to express unknown/unexpressed quantities.

OpenStudy (lgbasallote):

hmm..so you have solved various algebra questions (even ones on discriminants) and you're asking algebra? heh :) anyway what do you need help on?

OpenStudy (lalaly):

@Mimi_x3 lol

OpenStudy (anonymous):

help me .....

Parth (parthkohli):

Unknown quantities are mostly expressed by the letters in English letters. \(x\) is an example.

OpenStudy (anonymous):

whats discriminant ..... it is an english language word ..

OpenStudy (anonymous):

right ??

Parth (parthkohli):

\(ax^2 + bx + c = 0 \Longrightarrow D = b^2 - 4ac\)

OpenStudy (anonymous):

@lalaly : exxplain please ///

OpenStudy (anonymous):

@Mimi_x3 : virus help :P

Parth (parthkohli):

Inexpressible terms are through the Greek letters(most) \(\imath\) \(\gamma\) \(\pi\)

OpenStudy (unklerhaukus):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}=\frac{-b\pm\sqrt{\Delta}}{2a}\] the discriminant is delta \[\Delta =b^2-4ac\]

OpenStudy (anonymous):

\[x=(-b \pm \sqrt{b^2-4ac})/2a\]

Parth (parthkohli):

See there? \(\Delta \) is another Greek letter.

OpenStudy (mimi_x3):

lol, seeking for my help? :P

OpenStudy (anonymous):

whats delta ,,,, psi ,,,,, pie ..... i dont know :'( ... i am even confused in differentiation ....i dont know that